# Applying the Viterbi Algorithm to Planetary-Mass Black Hole Searches

### v1

@@ I Introduction
The structure of the paper is the following: In Sec. II, we first present a summary of the Viterbi algorithm and discuss its theoretical basis. Subsequently, we show that the length of the SFT needs to be optimized based on the chirp mass of the binary system. Then, we provide a detailed description of how to compute the optimum length. Following that, in Sec. III, we demonstrate the application of SOAP package to the search of planetary mass binary black holes by preparing mock data containing Gaussian noise and injections of signals, with varying luminosity distances and chirp masses. Finally, in Sec. IV, we present the conclusions of our study and the outlook for future investigations. Appendix A demonstrates that higher-order PNs are indeed non-negligible for most of the parameter space, supporting the advantage of model-agnostic nature of our method.
@@ II.1 Viterbi Algorithm Framework
where we have used the fact that we can factorize the marginalized likelihood $p(D\mid{\bm{\nu}})$ as
@@
$\begin{split}\log p(\hat{\bm{\nu}}|D)=\max_{{\bm{\nu}}}{\biggl{\{}\log p(\nu_{0})+\log p({\bm{x}_{0}}|\nu_{0})}\\ \left.\sum_{i=1}^{N-1}\biggl{[}\log p(\nu_{i}|\nu_{i-1})+\log p({\bm{x}_{i}}|\nu_{i})\biggr{]}\right\}+\text{const.}\,,\end{split}$ (5)
@@ II.2 Short-Time Fourier Transforms
where $N_{f}$ is the number of discrete frequency channels and $k$ is the label for GW frequency bins. Here, the variable $m$ is a dummy time index, and $n$ pinpoints the location of the window along the time axis.
@@
$T_{\mathrm{SFT}}\geq t_{\mathrm{GW}}(f+\frac{1}{2T_{\mathrm{SFT}}})-t_{\mathrm{GW}}(f-\frac{1}{2T_{\mathrm{SFT}}})\,,$ (9)
@@ II.3 Optimum SFT length
where tilde denotes the Fourier transform. The optimum SNR $\rho^{\mathrm{opt}}_{i}$ is given by
@@
For very large $N_{\mathrm{SFT}}$, the higher order moments tend to 0, and as expected from the central limit theorem, the distribution of Eq. (18) can be approximated by a Gaussian with the mean and standard deviation given in Eq. (II.3). Assuming that $N_{\mathrm{SFT}}\gg 1$ and a Gaussian is a good approximation, the number of “sigmas” $n_{\sigma}$ at which a signal with optimum SNR $\rho^{\mathrm{opt}}_{\mathrm{tot}}$ is expected to be is given by
@@
Figure 2: Colormap of $n_{\sigma}$ given by Eq. (21), normalized to the maximum possible $n_{\sigma}$ that can be obtained at each chirp mass as a function of the SFT length and the chirp mass. We also show with a solid black line the optimum SFT length as a function of the chirp mass. We assume the same typical LIGO O3 PSD of Fig. 1, and use the optimum low-frequency cutoff of $f_{0}=61.1\mathrm{Hz}$.
@@
Note that this function is solely determined by the shape of the PSD and does not depend on the signal parameters, such as the chirp mass. In Fig. 1, we show the colormap of $F(f_{0},f_{*})$ computed for a typical LIGO O3 PSD Abbott et al. (2020), where we observe that there is a well-defined maximum at $f_{0,\mathrm{opt}}=61.1$Hz and $f_{*,\mathrm{opt}}=126.8$Hz. To obtain this value of $f_{*}$, the length of the optimum SFT length depends on the chirp mass as given in Eq. (11)
@@
$d_{L}^{\mathrm{coh}}\!=\!0.544~{}\mathrm{Mpc}\!\!\left(\!\frac{\mathcal{M}_{c}}{10^{-2}M_{\odot}}\!\right)^{\!5/6}\!\!\!\left(\!\frac{n_{\sigma}}{10}\!\right)^{\!-1}\!\!\!\left(\!\frac{\int_{f_{\mathrm{min}}}^{f_{\mathrm{max}}}\frac{df}{f^{7/3}S_{n}(f)}}{3.85\times 10^{43}\mathrm{Hz}^{-1/3}}\!\right)^{\!\!1/2}.$ (34)
@@ II.4 Constraint on PBH abundance
Figure 3: Left panel: The Root MeanSquare (RMS) differences between the detected signal by the algorithm and the zero-noise injection as a function of the luminosity distances ($d_{L}$) of the sources. The RMS is given in units of frequency bins. Right panel: The Root MeanSquare (RMS) differences between the detected signal by the algorithm and the zero-noise injection as a function of the matched filter SNR for various cases of luminosity distances ($d_{L}$).
Figure 5: The normalized maximum probability of signal detection as a function of the luminosity distance for the cases of $\mathcal{M}_{c}=10^{-2}M_{\odot}$ and $\mathcal{M}_{c}=5\times 10^{-3}M_{\odot}$.
Figure 6: The average amount of time needed in order to complete signal detection with a standard M1 MacBook Pro for $\mathcal{M}_{c}=10^{-2}M_{\odot}$.

### v3

@@ I Introduction
The structure of the paper is the following: In Sec. II, we first present a summary of the Viterbi algorithm and discuss its theoretical basis. Subsequently, we show that the length of the SFT needs to be optimized based on the chirp mass of the binary system. Then, we provide a detailed description of how to compute the optimum length. We also provide an estimation of the sensitivity in terms of distance and its implications for the PBH abundance that could be obtained through this search. Additionally, we discuss the detection statistics with a detailed investigation into their distribution. Following that, in Sec. III, we demonstrate the application of SOAP package to the search of planetary mass binary black holes by preparing mock data containing Gaussian noise and injections of signals, with varying luminosity distances and chirp masses. Finally, in Sec. IV, we present the conclusions of our study and the outlook for future investigations. Appendix A demonstrates that higher-order PNs are indeed non-negligible for most of the parameter space, supporting the advantage of the model-agnostic nature of our method.
@@ II.1 Viterbi Algorithm Framework
where we have used the fact that we can factorize the likelihood $p(D\mid{\bm{\nu}})$ as
@@
$\begin{split}\log p(\hat{\bm{\nu}}|D)=\max_{{\bm{\nu}}}{\biggl{\{}\log p(\nu_{0})+\log p({\bm{x}_{0}}|\nu_{0})}\\ \left.+\sum_{i=1}^{N-1}\biggl{[}\log p(\nu_{i}|\nu_{i-1})+\log p({\bm{x}_{i}}|\nu_{i})\biggr{]}\right\}+\text{const.}\,,\end{split}$ (5)
@@ II.2 Short-Time Fourier Transforms
where $N_{f}=T_{\mathrm{SFT}}f_{\mathrm{s}}/2$, with $f_{\mathrm{s}}$ being the sampling frequency, is the number of discrete frequency channels and $k$ is the label for GW frequency bins. Here, the variable $m$ is a dummy time index, and $n$ pinpoints the location of the window along the time axis.
@@
$T_{\mathrm{SFT}}\leq t_{\mathrm{GW}}(f+\frac{1}{2T_{\mathrm{SFT}}})-t_{\mathrm{GW}}(f-\frac{1}{2T_{\mathrm{SFT}}})\,,$ (9)
@@ II.3 Optimum SFT length
where the tilde denotes the Fourier transform. The optimum SNR $\rho^{\mathrm{opt}}_{i}$ is given by
@@
For very large $N_{\mathrm{SFT}}$, the higher order moments tend to 0, and as expected from the central limit theorem, the distribution of Eq. (18) can be approximated by a Gaussian with the mean and standard deviation given in Eq. (20). Assuming that $N_{\mathrm{SFT}}\gg 1$ and a Gaussian is a good approximation, the number of “sigmas” $n_{\sigma}$ at which a signal with optimum SNR $\rho^{\mathrm{opt}}_{\mathrm{tot}}$ is expected to be is given by
@@
Figure 2: Colormap of $n_{\sigma}$ given by Eq. (21), normalized to the maximum possible $n_{\sigma}$ that can be obtained at each chirp mass as a function of the SFT length and the chirp mass. We also show with a solid black line the optimum SFT length as a function of the chirp mass. We assume the same typical LIGO O3 PSD of Fig. 1 and use the optimum low-frequency cutoff of $f_{0}=61.1\mathrm{Hz}$.
@@
Note that this function is solely determined by the shape of the PSD and does not depend on the signal parameters, such as the chirp mass. In Fig. 1, we show the colormap of $F(f_{0},f_{*})$ computed for a typical LIGO O3 PSD Abbott et al. (2020), where we observe that there is a well-defined maximum at $f_{0,\mathrm{opt}}=61.1$Hz and $f_{*,\mathrm{opt}}=126.8$Hz. To obtain this value of $f_{*}$, the optimum SFT length depends on the chirp mass as given in Eq. (11)
@@
$d_{L}^{\mathrm{coh}}\!=\!0.544\mathrm{Mpc}\!\!\left(\!\frac{\mathcal{M}_{c}}{10^{-2}M_{\odot}}\!\right)^{\!\!5/6}\!\!\!\left(\!\frac{n_{\sigma}}{10}\!\right)^{\!\!-1}\!\!\!\left(\!\frac{\int_{f_{\mathrm{min}}}^{f_{\mathrm{max}}}\frac{df}{f^{7/3}S_{n}(f)}}{3.85\times 10^{43}\mathrm{Hz}^{-1/3}}\!\right)^{\!\!\!1/2}.$ (34)
@@
Comparing the predicted sight distances of the coherent (Eq. (33)) and incoherent (Eq. (29)) cases, it might appear that we could have $d_{L}^{\mathrm{coh}}<d_{L}^{\mathrm{inc}}$, contradicting the expectation that a fully-coherent matched-filter search is optimal. Indeed, the ratio between both sight distances is given by
$\frac{d_{L}^{\mathrm{inc}}}{d_{L}^{\mathrm{coh}}}=\frac{n_{\sigma}^{\mathrm{coh}}}{\sqrt{2n_{\sigma}^{\mathrm{inc}}\sqrt{N_{\mathrm{SFT}}}}}\,,$ (35)
where we have assumed that the coherent and incoherent searches cover the same frequency range, i.e. $f_{\mathrm{min}}=f_{0}$ and $f_{\mathrm{max}}=f_{*}$ in Eq. (33), and we have used
$N_{\mathrm{SFT}}=\frac{T_{\mathrm{CBC}}}{T_{\mathrm{SFT}}}=\frac{\sqrt{30}}{64\pi^{4/3}}f_{*}^{11/6}f_{0}^{-8/3}\left(\frac{G\mathcal{M}_{c}}{c^{3}}\right)^{-5/6}\,.$ (36)
A priori, one could think that the ratio $d_{L}^{\mathrm{inc}}/d_{L}^{\mathrm{coh}}$ of Eq. (35) can become greater than 1, if $n_{\sigma}^{\mathrm{coh}}>\sqrt{2n_{\sigma}^{\mathrm{inc}}\sqrt{N_{\mathrm{SFT}}}}$. However, we have to take into account that $n_{\sigma}^{\mathrm{coh}}$ and $n_{\sigma}^{\mathrm{inc}}$ indicate different confidences due to the difference in their distributions, and we should compare them at a fixed confidence. In the case of the coherent search, $n_{\sigma}^{\mathrm{coh}}$ follows a normal $\mathcal{N}(0,1)$ distribution, and therefore the false alarm probability (FAP) is
$\mathrm{FAP}_{\mathrm{coh}}=P_{\mathcal{N}(0,1)}(\rho^{\mathrm{mf}}>n_{\sigma}^{\mathrm{coh}})=\frac{1}{2}\mathrm{erfc}\left(\frac{n_{\sigma}^{\mathrm{coh}}}{\sqrt{2}}\right)\,,$ (37)
where $\mathrm{erfc}(x)$ is the complementary error function. In the incoherent case,$n_{\sigma}^{\mathrm{inc}}$ follows a $\chi^{2}$ distribution with $2N_{\mathrm{SFT}}$ degrees of freedom, and therefore:
$\mathrm{FAP}_{\mathrm{inc}}$ $=P_{\chi^{2}_{2N_{\mathrm{SFT}}}}(\rho_{\mathrm{tot}}^{2}>2N_{\mathrm{SFT}}+2\sqrt{N_{\mathrm{SFT}}}n_{\sigma}^{\mathrm{inc}})$
$=Q(N_{\mathrm{SFT}},N_{\mathrm{SFT}}+\sqrt{N_{\mathrm{SFT}}}n_{\sigma}^{\mathrm{inc}})\,,$ (38)
where $Q(a,x)$ is the regularized upper incomplete gamma function, defined as
[…]
When both methods have the same FAP, the values of $n_{\sigma}$ for the coherent and incoherent searches are related by
$\frac{1}{2}\mathrm{erfc}\left(\frac{n_{\sigma}^{\mathrm{coh}}}{\sqrt{2}}\right)=Q(N_{\mathrm{SFT}},N_{\mathrm{SFT}}+\sqrt{N_{\mathrm{SFT}}}n_{\sigma}^{\mathrm{inc}})\,.$ (40)
In Fig. 3, we plot the ratio $d_{L}^{\mathrm{inc}}/d_{L}^{\mathrm{coh}}$ by substituting this relation to Eq. (35). We observe that $d_{L}^{\mathrm{inc}}/d_{L}^{\mathrm{coh}}<1$ for all plotted values of $N_{\mathrm{SFT}}$ and $n_{\sigma}^{\mathrm{coh}}$, confirming that the coherent search is always more sensitive than the incoherent one. We also plot the line $n_{\sigma}^{\mathrm{coh}}=\sqrt{N_{\mathrm{SFT}}-1}$, which separates the two different behaviors of the ratio $d_{L}^{\mathrm{inc}}/d_{L}^{\mathrm{coh}}$. When $n_{\sigma}^{\mathrm{coh}}\ll\sqrt{N_{\mathrm{SFT}}-1}$, the $\chi^{2}_{2N_{\mathrm{SFT}}}$ distribution can be well approximated by a Gaussian, and therefore we expect $n_{\sigma}^{\mathrm{inc}}\approx n_{\sigma}^{\mathrm{coh}}$, and Eq. (35) reduces to
$\frac{d_{L}^{\mathrm{inc}}}{d_{L}^{\mathrm{coh}}}\approx\sqrt{\frac{n_{\sigma}^{\mathrm{coh}}}{2\sqrt{N_{\mathrm{SFT}}}}}.$ (41)
Therefore, in the limit of having a very large number of SFTs, the sight distance of the incoherent search is degraded proportional to $N_{\mathrm{SFT}}^{-1/4}$ compared to the coherent case. On the opposite limit, when $n_{\sigma}^{\mathrm{coh}}\gg\sqrt{N_{\mathrm{SFT}}-1}$, we can approximate the ratio $d_{L}^{\mathrm{inc}}/d_{L}^{\mathrm{coh}}$ of Eq. (35) as
$\frac{d_{L}^{\mathrm{inc}}}{d_{L}^{\mathrm{coh}}}\approx 1-\frac{2N_{\mathrm{SFT}}-1}{2(n_{\sigma}^{\mathrm{coh}})^{2}}\log\!\left(\!1+\frac{(n_{\sigma}^{\mathrm{coh}})^{2}}{2N_{\mathrm{SFT}}}\!\right)\,,$ (42)
which again confirms that the ratio is always smaller than 1. In Eq. (42) we also observe that $d_{L}^{\mathrm{inc}}/d_{L}^{\mathrm{coh}}\approx 1$ when $N_{\mathrm{SFT}}\ll(n_{\sigma}^{\mathrm{coh}})^{2}$, indicating that in this regime the sensitivity lost by using an incoherent search is small.
Figure 3: Ratio between the coherent and incoherent sight distances computed, using Eq. (35) and applying the condition of Eq. (40) to impose both to correspond to the same FAP. The contour lines of the ratio are shown as a function of the number of sigmas of the coherent search $n_{\sigma}^{\mathrm{coh}}$ and the number of SFTs $N_{\mathrm{SFT}}$ (or equivalently the chirp mass $\mathcal{M}_{c}$, which is related via Eq. (36)) and is shown on the upper horizontal axis. The red curve is plotted to show the boundary of $n_{\sigma}^{\mathrm{coh}}=\sqrt{N_{\mathrm{SFT}}-1}$, below which the distribution of $n_{\sigma}^{\mathrm{inc}}$ can be approximated by a Gaussian.
@@ II.4 Constraint on PBH abundance
II.5 Distribution of the Viterbi detection statistic
In this subsection, we discuss the distribution of the total SNR of the track recovered by the Viterbi algorithm in detail. The distribution of the total SNR squared $\rho_{\mathrm{tot}}^{2}$, defined in Eq.(12), is described by Eq.(18) if the track is arbitrarily chosen. However, to be precise, this is not the case because the Viterbi algorithm selects the track with the maximum value of $\rho_{\mathrm{tot}}^{2}$ out of all possible tracks allowed by a given transition matrix. Consequently, $\max_{\vec{\nu}}\rho^{2}_{\mathrm{tot},\vec{\nu}}\equiv\rho^{2}_{\mathrm{tot},\mathrm{max}}$ does not follow the $\chi^{2}$ distribution of Eq. (18).
Given a flat transition matrix that enables to jump to $n_{J}$ frequency segments away per time step (where $n_{J}=1$ is the setting in our case), the probability to calculate Eq. (2) is given by
$p(\nu_{i}|\nu_{i-1})=\begin{cases}\frac{1}{2n_{J}+1}&\;\frac{1}{\Delta f}|\nu_{i}-\nu_{i-1}|\leq n_{J}\\ 0&\;\frac{1}{\Delta f}|\nu_{i}-\nu_{i-1}|>n_{J}\\ \end{cases}\,.$ (47)
where $\Delta f=1/T_{\mathrm{SFT}}$ is the frequency resolution of the SFT. There are $12n_{J}+$ 1 possible ways at each time step, and this is raised to the power of $N_{\rm SFT}$ when considering combinations across all time steps. Then, as the algorithm equally explores all initial frequencies, the total number of tracks that the Viterbi algorithm maximizes over is obtained by multiplying the number of frequency segments $n_{f}=T_{\mathrm{SFT}}(f_{\mathrm{max}}-f_{\mathrm{min}})$,
$N_{t}=n_{f}(2n_{J}+1)^{N_{\mathrm{SFT}}}\,.$ (48)
For typical values of $n_{J}=1$ and $N_{\mathrm{SFT}}\sim 1000$, the number of possible tracks becomes huge, of order $\sim 10^{500}$.
If we assume the tracks to be independent, computing the distribution of the maximum total SNR squared among these tracks is simple (see Appendix B). However, the correlations between tracks are crucial. Given that different SFT segments are uncorrelated, we have
$\mathrm{Cov}(|\rho^{\mathrm{mf}}_{ij}|^{2},|\rho^{\mathrm{mf}}_{kl}|^{2})=4\delta_{ij,kl}\,,$ (49)
where $\rho^{\mathrm{mf}}_{ij}$ denotes the matched filter SNR in the $j$-th frequency of the $i$-th SFT segment, which are distributed as uncorrelated complex Gaussians. Using Eq. (49), we find that the correlation between two different tracks is given by
$r(\rho^{2}_{\mathrm{tot},\vec{\nu}_{1}},\rho^{2}_{\mathrm{tot},\vec{\nu}_{2}})\equiv\frac{\mathrm{Cov}(\rho^{2}_{\mathrm{tot},\vec{\nu}_{1}},\rho^{2}_{\mathrm{tot},\vec{\nu}_{2}})}{\sqrt{\mathrm{Var}(\rho^{2}_{\mathrm{tot},\vec{\nu}_{1}})\mathrm{Var}(\rho^{2}_{\mathrm{tot},\vec{\nu}_{2}})}}=\frac{n_{c}}{N_{\mathrm{SFT}}}\,,$ (50)
where $n_{c}$ is the number of segments that tracks have in common.
Given the complexity of the problem, in practice, we investigate the distribution of $\rho^{2}_{\mathrm{tot},\mathrm{max}}$ using simulated background data assuming Gaussian noise. In Fig. 4, we plot the mean and standard deviation of this distribution as a function of the number of SFTs, $N_{\mathrm{SFT}}$ considered in the analysis, while fixing $n_{f}=1000$ and $n_{J}=1$, which are the values consistent with our mock data analysis in the subsequent section. In the simulations, we observe that the distribution depends very weakly on $n_{f}$, similarly to what happens in the uncorrelated case discussed in Appendix B. This can be expected writing Eq. (48) as $N_{t}=(2n_{J}+1)^{N_{\mathrm{SFT}}+\log(n_{f})}$, and observing that usually $\log(n_{f})\ll N_{\mathrm{SFT}}$.
Figure 4: Mean $\mu$ and standard deviation $\sigma$ of the distribution of the maximum over Viterbi tracks of the total SNR $\rho^{2}_{\mathrm{tot},\mathrm{max}}$ as a function of the number of SFTs, $N_{\mathrm{SFT}}$. Note that the vertical axis has different scales for the mean (left) and the standard deviation (right). We assume $n_{f}=1000$ and $n_{J}=1$, consistent with the values used in our analysis of Sec. III.
In Fig. 4, we find that, similarly to Eq. (20), in the absence of signal ($\rho^{\mathrm{opt}}_{\mathrm{tot}}=0$), the mean is proportional to $N_{\mathrm{SFT}}$ and the standard deviation is proportional to $\sqrt{N_{\mathrm{SFT}}}$. The proportionality constants are, however, very different, and we find
$\mu(\rho^{\mathrm{opt}}_{\mathrm{tot}}=0,N_{\mathrm{SFT}})$ $=4.8N_{\mathrm{SFT}}\,,$ (51a)
$\sigma(\rho^{\mathrm{opt}}_{\mathrm{tot}}=0,N_{\mathrm{SFT}})$ $=0.74\sqrt{N_{\mathrm{SFT}}}\,.\;$ (51b)
In Fig. 5, we show how the probability distribution function (PDF) of $\rho^{2}_{\mathrm{tot},\mathrm{max}}$ varies as a function of $N_{\mathrm{SFT}}$. We can observe that for $N_{\mathrm{SFT}}=1$, it resembles a Gumbel distribution, as expected from Appendix B, since in this case there is no correlation between paths. However, as $N_{\mathrm{SFT}}$ increases, the distribution becomes more symmetric, resembling something more similar to a Generalized Extreme Value (GEV) distribution
$P_{\mathrm{GEV}}(x)=\exp\left\{-(1+\gamma x)^{-1/\gamma}\right\}\,,$ (52)
which approaches the Gumbel distribution in the limit of $\gamma\to 0$. Fitting the GEV distribution to the $N_{\mathrm{SFT}}=1000$ data, we obtain a best fit value of $\gamma=-0.133$, shown also in Fig. 5.
Figure 5: PDF of $\rho^{2}_{\mathrm{tot},\mathrm{max}}$ scaled and shifted to have zero mean and a standard deviation of 1.
Similarly to what is done with the critical ratio (CR) of the Hough transform Miller et al. (2018), to preselect events when using the Viterbi algorithm, we can look at the number of standard deviations from the mean the observed total SNR of the reconstructed track is at
$n_{\sigma}=\frac{\rho^{2}_{\mathrm{tot},\mathrm{max}}-\mu(\rho^{\mathrm{opt}}_{\mathrm{tot}}=0,N_{\mathrm{SFT}})}{\sigma(\rho^{\mathrm{opt}}_{\mathrm{tot}}=0,N_{\mathrm{SFT}})}\,.$ (53)
To evaluate the expectation value of $n_{\sigma}$, we can use the fact that, for a sufficiently loud signal, the Viterbi track corresponds to the true track, and in this case, $\mu(\rho^{\mathrm{opt}}_{\mathrm{tot}},N_{\mathrm{SFT}})$ is given by Eq. (20a). Therefore, we obtain
$\langle n_{\sigma}\rangle$ $\approx\frac{2N_{\mathrm{SFT}}+(\rho^{\mathrm{opt}}_{\mathrm{tot}})^{2}-\mu(\rho^{\mathrm{opt}}_{\mathrm{tot}}=0,N_{\mathrm{SFT}})}{\sigma(\rho^{\mathrm{opt}}_{\mathrm{tot}}=0,N_{\mathrm{SFT}})}\,.$ (54)
Substituting the approximations of $\mu$ and $\sigma$ shown in Eq. (51), for the $n_{J}=1$ case, $\langle n_{\sigma}\rangle$ is approximately given by
$\langle n_{\sigma}\rangle$ $\approx\frac{(\rho^{\mathrm{opt}}_{\mathrm{tot}})^{2}-2.88N_{\mathrm{SFT}}}{0.74\sqrt{N_{\mathrm{SFT}}}}\,.$ (55)
We expect that, in the absence of signal $(\rho^{\mathrm{opt}}_{\mathrm{tot}})^{2}\to 0$, we would have $\langle n_{\sigma}\rangle\to 0$. However this is not the behaviour we observe in Eq. (54) because of the loud signal assumption. We see $\langle n_{\sigma}\rangle$ vanishes when $(\rho^{\mathrm{opt}}_{\mathrm{tot}})^{2}=\mu(\rho^{\mathrm{opt}}_{\mathrm{tot}}=0,N_{\mathrm{SFT}})-2N_{\mathrm{SFT}}$, and we will see in the next section that, at this point, Eq. (54) no longer serves as a good approximation. The approximation is broken because, as the signal becomes fainter, we do not expect the Viterbi algorithm to recover the true track but a noise-dominated track.
Note that in the derivation of the optimum length of the SFT, done in Sec. II.3, we did not consider the trials factor of the Viterbi algorithm, and therefore the form of $n_{\sigma}$ in Eq.(21), which is used to optimize the SFT length, is much simpler. However, in the large SNR limit in which a signal can be detected and where Eq. (55) applies, we do not expect the trials factor to be so relevant, since we find the same dependence of $n_{\sigma}\propto(\rho^{\mathrm{opt}}_{\mathrm{tot}})^{2}/\sqrt{N_{\mathrm{SFT}}}$. Therefore, even if the correction by the trials factor may introduce a factor of difference in the sensitivity estimate, we expect the estimates of Sec. II.3, in which we have full analytical control, to be broadly applicable.
Figure 6: Left panel: The Root Mean Square (RMS) differences between the detected signal by the algorithm and the zero-noise injection as a function of the luminosity distances ($d_{L}$) of the sources. The RMS is given in units of frequency bins. Right panel: The RMS differences between the detected signal by the algorithm and the zero-noise injection as a function of the total SNR for various cases of luminosity distances ($d_{L}$).